OpenStudy (juscallmesteve):

Linearization of F(x), x= a f(x) = 1/(6x+2) a=0

OpenStudy (juscallmesteve):

so i take the derivative and f(x) i get \[\frac{ 6 }{ (6x+2)^2 }\]

OpenStudy (juscallmesteve):

plug 0 into both f(x) and \[f \prime (x) \]

OpenStudy (juscallmesteve):

f(0) = 1/2 and f'(0) = -1.5 so plug into linearization formula \[L(x) = F(0) + F'(0)(x-n)\] this comes out to L(x) = 1/2 + (-1.5)(x-0) is that correct?

OpenStudy (anonymous):

your derivative is wrong.

OpenStudy (juscallmesteve):

forgot the (-)

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

oh, you solved using the negative but forgot to write it above yeah you're right